OpenStudy (anonymous):

The legs of an isosceles triangle have lengths x + 1 and -x + 7 . The base has length 3x - 3. What is the length of the base? A. 4 B. 6 C. 3 D. cannot be determined

OpenStudy (alekos):

The legs of an isosceles triangle should have the same length by definition

OpenStudy (anonymous):

first, set it up so that: x+1=-x+7 then add x to both sides: 2x+1=7 subtract 1 from both sides: 2x=6 solve for x by dividing both sides by 2: x=3

OpenStudy (anonymous):

since we know that x=3, substitute 3 in for 3 for the base